OpenStudy (anonymous):

simplify completely ((x2-4x+4)/(x2+10x+25)) x ((x+5)/(x2+3x-10))

OpenStudy (anonymous):



OpenStudy (diyadiya):

Factorise (x2-4x+4) & (x2+10x+25)

OpenStudy (anonymous):

Diya is smart you can see ^^

OpenStudy (diyadiya):

\[\huge LoL\]

OpenStudy (anonymous):

That's huge lol :D

OpenStudy (diyadiya):

\[\huge x^2-4x+4 =(x-2)(x-2)\]

OpenStudy (anonymous):

You should get \( \huge \frac{(x-2)}{(5+x)^2} \) as final answer

OpenStudy (anonymous):

thank you i was simplifying it but im not sure about my signs, Fool your a big help...ty

OpenStudy (anonymous):

so are you Diya

OpenStudy (anonymous):

Thanks, but all my answers are pure guesses ;)

OpenStudy (anonymous):

WHAT?!?!?!

OpenStudy (diyadiya):

\[\huge x^2-4x+4 =(x-2)(x-2)\]\[\huge x^2+10x+25=(x+5)(x+5)\] \[\huge \frac{ (x-2)^2}{(x+5)^2} \times \frac{(x+5)}{(x-2)(x+5)}\] \[\huge \frac{(x-2)}{(x+5)^2}\]

OpenStudy (diyadiya):

you're welcome!